# Complexity at Scale: A Quantitative Analysis of an Alibaba Microservice Deployment

### v2

@@ 4.1. Microservice Instances
MSs at Alibaba and Meta demonstrate similar distributions of, and maximum, instance counts: the deployed instances per MS at both MSAa demonstrate a heavy-tail (see Table 2), however, the median, average and 99-th percentile values for Meta are an order of magnitude higher. Given that the number of instances deployed at Meta is also an order of magnitude higher (see Section 3.1), this means that the central tendencies for both distributions when normalising by the total number of instances are similar, and thus the typical MS within both deployments represents a similar fraction of the overall deployed instances. Notably, we found that the maximum number of instances for a single MS were remarkably similar (227,000 vs 557,000) despite this order of magnitude difference in the total number of instances. This indicates that the heavy tail is more pronounced in this dataset, with the MS with the most instances accounting for 12.4% of the total number of instances compared to Meta’s 4.6%. The convergence of maximum values despite the difference in the total number of instances might suggest that there are common operational ceilings when it comes to scaling. This ceiling could be a result of coordination overhead, such as managing load balancing between a larger number of instances, or non-linear performance gains due to resource contention for shared resources such as databases, caches, and network bandwidth.
@@ 4.2. Nodes and Microservices
MS instances can be deployed across a large number of host machines: as MS instances may be transient, we analysed the MS instances and their associated nodes within the first minute of the resource data. This time period was arbitrarily chosen, but assumed to be representative of the dataset. Our analysis revealed that 84% of these MSs were concurrently deployed on more than one node - with the median being 4 nodes and one MS spanning as many as 4,114 nodes (see Figure 4(a)). Moreover, we found that 96% of MSs had a single instance per node, as shown in Figure 4(b). These findings are likely a result of a deliberate orchestration strategy to aid load balancing and reliability by spreading MSs across nodes. As found in our results, in a minority of cases, MSs may be deployed multiple times on the same node to reduce fragmentation.
@@ 4.3. Microservice Workload
The distribution of workload across MSs is extremely uneven: the accumulation of calls from upstream MSs over a period of time can be defined as MS workload (Luo et al., 2022). To investigate how this workload was distributed across MSs, we aggregated the total calls received by each MS over the full 14-day period. Our analysis revealed a vast variation in workload across MSs, receiving from a single call to as many as 6.9 billion calls, the latter representing 7.1% of all calls made over the two weeks (see Figure 5). The cumulative workload distribution was heavy-tailed, with 21% of MSs receiving fewer than 10 calls, whereas the top 1% accounted for 87% of all calls (these workloads are likely associated with e-commerce application traffic which generates the most workload for the Alibaba deployment). These results corroborate previous research reporting heavy-tailed MS CPU workloads at Alibaba (Luo et al., 2022).
Figure 6. Cumulative frequency of the aggregate in-degree (blue) and out-degree (pink) per MS, with lower and upper quartiles (dash–dot) and median (dashed).
4.4. Microservice Dependencies
A complementary perspective to the MSA’s global dependency structure explored in Section 3.4, comes from analysing the local structure, such as MS in-degree and out-degree (Luo et al., 2021; Huye et al., 2023). Here, the in-degree of a MS indicates how many other MSs call it, while the out-degree reflects how many other MSs it calls. We argue that these more aptly capture the importance of a MS within the architecture. A high in-degree signals a critical hub that many MSs rely on and is shared across a large amount of front-end functionality, and a high out-degree suggests a highly composite MS that aggregates the functionality of many downstream MSs.
A small number of MSs have a disproportionately large number of dependencies: using the aggregate dependency network constructed in Section 3.4, we found that both in-degree and out-degree distributions exhibited a heavy-tail (see Figure 6). Of the MSs, 38.5% had an in-degree of 1, with a median of 2, yet the maximum in-degree reached 13,030, with one MS being called by 19.2% of all other MSs. In contrast, only 20.2% of MSs had an out-degree of 1, and the median out-degree was 4. However, the maximum out-degree was lower at 2,245 - meaning that 3.31% of all MSs are called by this MS. Despite the higher maximum value for in-degree, these distributions indicate that generally most MSs call more peers than they are called by, in contrast to the findings of Meta (Huye et al., 2023), but supported by previous research at Alibaba (Luo et al., 2021). Similarly to workload, many MSs with high numbers of dependencies are likely associated with e-commerce applications. However, it is also likely that some also represent common core MSs that are shared between applications and business domains (Newman, 2021) which may be central to overall system functionality; these architectural hot spots have also been identified in previous research (Luo et al., 2021).
(a) Kendall’s Tau correlation
(b) Jaccard similarity for top 1%
Table 3. Degree Comparison Between Alibaba and Meta
@@ 4.4. Microservice Dependencies
In‑Degree Out‑Degree In‑Degree Out‑Degree
@@
Alibaba and Meta’s local dependency topology structure is similar: in Section 3.4 we noted that this dataset’s global dependency structure was considerably sparser than that of Meta’s. Yet, we found the local degree distributions to be strikingly similar, as shown in Table 3. Greater sparsity, in conjunction with a larger number of MSs, and a local structure resembling Meta’s, indicates that new MSs are incorporated by establishing dependencies with only a limited subset of peers. This observation, coupled with the similarity in local dependency structures, that likely arise from different organisational structures, technology stacks, and scales at Alibaba and Meta, suggests a convergence in MSA design. Such convergence likely arises due to universal constraints, such as performance, reliability, and maintainability, driving MSA design toward bounded complexity.
@@ 4.5. Instances, Workload, and Dependencies
There is only a weak relationship between the number of MS instances and dependencies, and their workload: so far we demonstrated that instances, workload, and dependencies all display heavy-tailed distributions. Here, we explore whether there exists a relationship between these metrics using Kendall’s Tau. The resultant correlation matrix shown in Figure 7(a) revealed a weak relationship between all studied variables. The highest correlation was found between total calls received and in-degree, suggesting that, to some extent, MSs that receive a wider array of different calls from more MSs also receive more calls overall; this provides some supporting evidence for the hub-like properties of some of these MSs discussed in Section 4.4. We only found a weak relationship between total calls received and number of instances deployed for a specific MS. This indicates that other factors such as workload variability, functionality CPU requirements, and MS-specific scaling rules, play an important role in shaping the number of deployed instances of a MS.
@@ 4.7. Call Graph Variability
Given the very large size of the dataset, we made several important design choices when reconstructing call graphs to reduce computational requirements. Firstly, we aggregated call graphs on a per-datafile basis, corresponding to 3 minutes of data in total. We found that a small fraction of call graphs (0.009%), mostly due to slow message queue calls averaging over 30 minutes, spanned multiple files. To handle these, we collected each affected Trace ID’s calls to the file of invocation within an 80‑file window, which we found covered 97% of all overlaps. Secondly, we did not reconstruct the full topological structure of call graphs in accordance with RPC IDs; instead we represented call graphs as de-duplicated sets of calls. To quantify the similarity between these sets, we then used LSH to generate hashes, which could then be used to efficiently estimate Jaccard similarity. Our approach of representing call graphs as de-duplicated and ordered vectors of MS call pairs has a few caveats for the interpretation of results. Firstly, invocation order is not preserved, meaning that two call graphs with the same calls that occur in different orders would be considered identical. However, the invocation order of MSs are not random as they are based on underlying logic, and therefore two call graphs that have identical MS invocations are likely to be very similar in topological structure. Secondly, the number of times a call occurs within a call graph is not considered in our set representation. This means that two identical call graphs may actually be different due to any number of repeated calls, meaning that the results presented below may underestimate true variability.
@@
One issue with attempting to quantify the variability of call graphs is that inconsistencies within the observability data can impact results. For example, call graphs can be incomplete due to randomly missing data, meaning that two call graphs could be determined to be different simply due to missing data in one or both call graphs, rather than due to a true difference in the actual call patterns. Indeed, the authors of (Huye et al., 2024) found that missing data is highly prevalent within the call data released by Alibaba. However, due to how prevalent this missing data is, it is likely due to a systematic lack of observability or specific choices made during data processing for publishing  [^3] Due to space limitations we have omitted our full analysis, but the evidence supports this conclusion.. Therefore, this missing data is deterministic and does not affect our call set representation comparisons.
@@
The number of unique call graphs is measured in millions and front-end functionality can be associated with a large number of different call graphs: it total we identified 79,737,924 unique call graph sets. However, the median number of unique call graph sets per service was only 1, with 59% of Service IDs being associated with a single call graph set, which is likely a product of the sporadic non-user facing functionality identified in Section 3.2. Nonetheless, a significant portion of Service IDs showed variability in their call graphs, with a mean number of unique call graph sets per service of 31, and a maximum number of unique call graph sets of just over 3 million.
@@
Call graphs from the same service are not necessarily similar in their MS calls: a large number of unique call graph sets does not necessarily imply a high degree of call graph variability per service, as each call graph may only display small deviations in call patterns. Therefore, we investigated the average Jaccard similarity between call graph sets for each front-end service that produced more than a single unique call graph set. As shown in Figure 9(a), the distribution of similarities, considering the frequency of each call graph occurrence shows a grouping of high average similarities between 0.65 and 0.9, accounting for 52.8% of all services. On the other hand, we also found that a substantial number of services (28.5%) demonstrated less than 0.5 similarity between their call graph sets on average, with 6.5% of service showing no similarity. As we account for the number of times each call graph set occurred in our analysis, the cluster of higher average similarities may be caused by dominant call graph sets per service. This follows from design principles, where a given front-end feature typically relies on a core set of MS calls with small deviations occurring due to conditional or alternative functionality.
@@
Front-end functionality often has a dominant call graph: to further understand the prevalence of dominant call graphs, we explored the proportion of the total call graph sets produced by each service that matched the most commonly occurring call graph. In a first analysis we found that there were a large number of services (13.4%) that only produced two unique call graphs each only appearing once, resulting in a spike of ratios at 0.5, these cases do not provide much information as to the prevalence of dominant call graphs and thus we excluded them from the subsequent analysis. The resulting distribution, shown in Figure 9(b), reveals a significant skew of the ratio of the dominant call graph set towards higher values than would be expected if all call graph sets had an equally likely probability to occurrence, with 82% of Service IDs demonstrating higher ratios than the maximum possible ratio if all call graphs were equally likely. This confirms that services generally have a dominant call graph that disproportionally occurs during service invocation, with rarer non-dominant call graphs.
@@
Non-dominant do not necessarily show high similarity to the dominant call graph pattern of a front-end service’s functionality: subsequently, we sought to characterise the level of similarity between dominant and non-dominant call graphs. We focused our investigation on services with a high prevalence of dominant call graphs (>>>0.75), for each Service ID we found the average Jaccard similarity between the dominant call graph set and all other call graph sets. The resulting distribution, shown in Figure 9(c), reveals that non-dominant call graphs show some similarity to non-dominant call graphs, with an average Jaccard similarity of 0.6. However, we found that 35% of non-dominant call graphs share less than 50% of the same calls, with 15% of non-dominant call graphs demonstrating zero similarity with corresponding dominant call graphs. Our findings extend previous observations of variability in MS call graphs within large-scale MSAs at both Alibaba and Meta (Luo et al., 2021; Huye et al., 2023). We found that not only are call graphs from the same front-end functionality or service potentially highly variable, but non-dominant call graphs can show dissimilarity in the types of MS calls involved. We note that the total number of unique call graphs is very large (79 million) even when not accounting for call graph structure or duplicate calls.

### v3

@@ 5.1. Nodes and Microservices
MS instances can be deployed across a large number of host machines: as MS instances may be transient, we analysed the MS instances and their associated nodes within the first minute of the resource data. This time period was arbitrarily chosen, but assumed to be representative of the dataset. Our analysis revealed that 84% of these MSs were concurrently deployed on more than one node - with the median being 4 nodes and one MS spanning as many as 4,114 nodes (see Figure 4(a)). Moreover, we found that 96% of MSs had a single instance per node, as shown in Figure 4(b). These findings are a result of a deliberate orchestration strategy to aid load balancing and reliability by spreading MSs across nodes (see 3). As found in our results, in a minority of cases, MSs may also be deployed multiple times on the same node to reduce fragmentation (see 3).
@@ 5.2. Microservice Workload
The distribution of workload across MSs is extremely uneven: the accumulation of calls from upstream MSs over a period of time can be defined as MS workload (luo_power_2022). To investigate how this workload was distributed across MSs, we aggregated the total calls received by each MS over the full 14-day period. Our analysis revealed a vast variation in workload across MSs, receiving from a single call to as many as 6.9 billion calls, the latter representing 7.1% of all calls made over the two weeks (see Figure 5). The cumulative workload distribution was heavy-tailed, with 21% of MSs receiving fewer than 10 calls, whereas the top 1% accounted for 87% of all calls. These workloads are likely to be associated with e-commerce application traffic, such as Taobao, which generates the most workload for the Alibaba deployment (see 3). These results corroborate previous research reporting heavy-tailed MS CPU workloads at Alibaba (luo_power_2022).
(a) CDF of aggregate in- and out-degree per MS (Alibaba).
@@
In-Degree Out-Degree In-Degree Out-Degree
@@
(b) Degree comparison between Alibaba and Meta.
Figure 6. Degree distribution and Alibaba–Meta comparison: (a) CDF of aggregate in-degree (blue) and out-degree (pink) per microservice at Alibaba, with lower/upper quartiles (dash–dot) and median (dashed); (b) summary statistics comparing in- and out-degrees across Alibaba and Meta.
5.3. Microservice Dependencies
A complementary perspective to the MSA’s global dependency structure explored in Section 4.4, comes from analysing the local structure, such as MS in-degree and out-degree (luo_characterizing_2021; huye_lifting_2023). Here, the in-degree of a MS indicates how many other MSs call it, while the out-degree reflects how many other MSs it calls. We argue that these more aptly capture the importance of a MS within the architecture. A high in-degree signals a critical hub that many MSs rely on and is shared across a large amount of front-end functionality, and a high out-degree suggests a highly composite MS that aggregates the functionality of many downstream MSs.
A small number of MSs have a disproportionately large number of dependencies: using the aggregate dependency network constructed in Section 4.4, we found that both in-degree and out-degree distributions exhibited a heavy-tail (see Figure 6(a)). Of the MSs, 38.5% had an in-degree of 1, with a median of 2, yet the maximum in-degree reached 13,030, with one MS being called by 19.2% of all other MSs. In contrast, only 20.2% of MSs had an out-degree of 1, and the median out-degree was 4. However, the maximum out-degree was lower at 2,245 - meaning that 3.31% of all MSs are called by this MS. Despite the higher maximum value for in-degree, these distributions indicate that generally most MSs call more peers than they are called by, in contrast to the findings of Meta (huye_lifting_2023), but supported by previous research at Alibaba (luo_characterizing_2021). Similarly to workload, many MSs with high numbers of dependencies are likely associated with e-commerce applications (see 3. However, it is also likely that some also represent common core MSs that are shared between applications and business domains (newman_building_2021) which may be central to overall system functionality; these architectural hot spots have also been identified in previous research (luo_characterizing_2021).
Alibaba and Meta’s local dependency topology structure is similar: in Section 4.4 we noted that this dataset’s global dependency structure was considerably sparser than that of Meta’s. Yet, we found the local degree distributions to be strikingly similar, as shown in Table 6(b). Greater sparsity, in conjunction with a larger number of MSs, and a local structure resembling Meta’s, indicates that new MSs are incorporated by establishing dependencies with only a limited subset of peers.
(a) Kendall’s Tau correlation
(b) Jaccard similarity for top 1%
This observation, coupled with the similarity in local dependency structures, that likely arise from different organisational structures, technology stacks, and scales at Alibaba and Meta, suggests a convergence in MSA design. Such convergence likely arises due to universal constraints, such as performance, reliability, and maintainability, driving MSA design toward bounded complexity.
@@ 5.4. Instances, Workload, and Dependencies
There is only a weak relationship between the number of MS instances and dependencies, and their workload: so far we demonstrated that instances, workload, and dependencies all display heavy-tailed distributions. Therefore, we explore whether there exists a relationship between these metrics using Kendall’s Tau (kendall_1938), a rank-based correlation measure that is less sensitive to extreme values than the conventional Pearson correlation coefficient. The resultant correlation matrix shown in Figure 7(a) revealed a weak relationship between all studied variables. The highest correlation was found between total calls received and in-degree, suggesting that, to some extent, MSs that receive a wider array of different calls from more MSs also receive more calls overall; this provides some supporting evidence for the hub-like properties of some of these MSs discussed in Section 5.3. We only found a weak relationship between total calls received and number of instances deployed for a specific MS. This indicates that other factors such as workload variability, functionality CPU requirements, and MS-specific scaling rules, play an important role in shaping the number of deployed instances of a MS.
@@ 5.6. Call Graph Variability
To efficiently evaluate and compare the diversity of call graphs within the Alibaba dataset we represented traces as vectors, a common approach for embedding traces (chen_tracemesh_2024). These vectors are then converted into lower-dimensional representations using locality-sensitive hashing via MinHash (broder_minhash_1997; indyk_lsh_1998). The pairwise similarity of the resulting hash signatures approximates the Jaccard similarity (broder_minhash_1997; indyk_lsh_1998; jaccard_similarity_1901), quantifying how similar traces are in terms of the types of calls they contain. Whilst this approach does not directly compare graph structure we argue that the invocation order of MSs is not random, being based on underlying logic, and therefore two call graphs that share identical MS invocations are likely to be very similar in topological structure. Nevertheless, given this caveat, the results presented below may underestimate the true variability of call graphs but still provide a reasonable approximation.
@@
A key challenge in quantifying the variability of call graphs is that inconsistencies in the observability data can impact results. For example, call graphs can be incomplete due to randomly missing data, meaning that two call graphs could be determined to be different simply due to missing data in one or both call graphs, rather than due to a true difference in the actual call patterns. Previous work has suggested that missing data is prevalent within the call data released by Alibaba. However, this missing data is due to a systematic lack of observability for specific MSs within the deployment, and types of calls not published in the trace data (see 3). Indeed, in a preliminary study of the first 12 hours of the dataset we found that only 0.12% of traces with missing data could not be accounted for by systematic observability loss. Therefore, because the vast majority of missing data is deterministic, it is unlikely to significantly bias our vector-based comparisons of call graphs.
@@
The number of unique call graphs is measured in millions and front-end functionalities can be associated with a large number of different call graphs: in total, we identified 79,737,924 unique call graph vectors. However, the median number of unique call graph vectors per service was only 1, with 59% of Service IDs being associated with a single call graph vector, which is likely a product of the sporadic non-user facing functionality identified in Section 4.2. Nonetheless, a significant portion of Service IDs showed variability in their call graphs, with a mean number of unique call graph vectors per service of 31, and a maximum number of unique call graph vectors of just over 3 million. These values are higher than found in previous work at Alibaba (luo_characterizing_2021), which may be explained by the use of graph learning and clustering, grouping together graphs that demonstrate similarity in their call graph structure.
@@
Call graphs from the same service are not necessarily similar in their MS calls: a large number of unique call graph vectors does not necessarily imply a high degree of call graph variability per service, as each call graph may only display small deviations in call patterns. Therefore, we investigated the average Jaccard similarity between call graph vectors for each front-end service that produced more than a single unique call graph vector. As shown in Figure 9(a), the distribution of similarities, considering the frequency of each call graph occurrence shows a grouping of high average similarities between 0.65 and 0.9, accounting for 52.8% of all services. On the other hand, we also found that a substantial number of services (28.5%) demonstrated less than 0.5 similarity between their call graph vectors on average, with 6.5% of service showing no similarity. As we account for the number of times each call graph vector occurred in our analysis, the cluster of higher average similarities may be caused by dominant call graph vectors per service. This follows from design principles, where a given front-end feature typically relies on a core set of MS calls with small deviations occurring due to conditional or alternative functionality.
@@
Front-end functionality often has a dominant call graph: to further understand the prevalence of dominant call graphs, we explored the proportion of the total call graph vectors produced by each service that matched the most commonly occurring call graph. In a first analysis we found that there were a large number of services (13.4%) that only produced two unique call graphs each only appearing once, resulting in a spike of ratios at 0.5, these cases do not provide much information as to the prevalence of dominant call graphs and thus we excluded them from the subsequent analysis. The resulting distribution, shown in Figure 9(b), reveals a significant skew of the ratio of the dominant call graph vector towards higher values than would be expected if all call graph vectors had an equally likely probability to occurrence, with 82% of Service IDs demonstrating higher ratios than the maximum possible ratio if all call graphs were equally likely. This confirms that services generally have a dominant call graph that disproportionally occurs during service invocation, with rarer non-dominant call graphs.
@@
Non-dominant do not necessarily show high similarity to the dominant call graph pattern of a front-end service’s functionality: subsequently, we sought to characterise the level of similarity between dominant and non-dominant call graphs. We focused our investigation on services with a high prevalence of dominant call graphs (>>0.75), for each Service ID we found the average Jaccard similarity between the dominant call graph vector and all other call graph vectors. The resulting distribution, shown in Figure 9(c), reveals that non-dominant call graphs show some similarity to non-dominant call graphs, with an average Jaccard similarity of 0.6. However, we found that 35% of non-dominant call graphs share less than 50% of the same calls, with 15% of non-dominant call graphs demonstrating zero similarity with corresponding dominant call graphs. Our findings extend previous observations of variability in MS call graphs within large-scale MSAs at both Alibaba and Meta (luo_characterizing_2021; huye_lifting_2023). We found that not only are call graphs from the same front-end functionality or service potentially highly variable, but non-dominant call graphs can show dissimilarity in the types of MS calls involved. We note that the total number of unique call graphs is very large (79 million) even when not accounting for call graph structure.
5.7. Heterogeneity Implications
MSs can be multiplexed by a large variety of other MSs and front-end functionality: a common assumption employed by many fault management approaches is that the behaviour of MSs is naturally correlated with the usage of front-end applications and their KPIs, and thus deviations in this relationship can be used to detect anomalies (samir_dla_2019; samir_dla_2019) and correlation with faulty KPIs can help localise the root cause of faults (wang_cloudranger_2018; ma_ms-rank_2019; ma_servicerank_2022; li_practical_2021; liu_microhecl_2021). However, we found that MSs can be involved with potentially thousands of front-end service functionalities (Section 5.5), inevitably resulting in signal mixing as the behaviour of MSs is the aggregate of many different front-end usage patterns. Without appropriate treatment of signal mixing, for instance through blind source separation (pal_blind_2013), these approaches are likely to incorrectly detect and attribute faults.
Another assumption employed by anomaly detection approaches is that metrics are approximately normally distributed and well behaved, allowing the use of the 3-sigma rule 3-Sigma rule (lin_microscope_2018; yu_microrank_2021; li_causal_2022; lee_eadro_2023) or the extreme studentized deviate test (vallis_novel_2014) to highlight outlier values. Instead, our findings demonstrate that MSs can be multiplexed by many front-end workloads, potentially producing non-Gaussian mixtures with varying means and variances, which can yield unexpected tail behaviours and non-stationary signals as different workloads dominate over time. These characteristics challenge the suitability of these statistical approaches in real-world systems which would experience increased false alarms and decreased accuracy under such conditions.
Another set of popular approaches for leveraging distributed tracing data for management is to treat calls within traces as sequences of events (nedelkoski_anomaly_2019; meng_midiag_2020; purfallah_mazraemolla_effective_2024). Sequence-learning approaches, such as recurrent neural networks, can then be used to learn the normal sequences of calls between MSs. Trained models can then be used to flag anomalous calls that fall below some threshold of expected next events. However, these approaches rely on the assumption that the number of next possible calls is moderate. However, we found that MSs can have very large out-degrees (Section 5.3), meaning a vast number of possible next events. This can dramatically increase the amount of training data required for such approaches. Additionally, a dispersed probability distribution over many potential next events could make approaches sensitive to threshold parameters. Furthermore, picking a one-size-fits-all threshold when the number of next events across MSs is heavy-tailed is likely to result in increased false negatives and increased false positives for MSs with low and high out-degree respectively.
Large‑scale MSAs generate many call graphs, with each front‑end function following a common structure but sometimes producing rare, divergent ones: an assumption of many management approaches is that normal call graphs are adequately represented within tracing data. This allows learning approaches to model the structure of normal call graphs (yang_transparently_2021; panahandeh_serviceanomaly_2024; wang_workflow-aware_2020; meng_detecting_2021) which can be subsequently used for anomaly detection and localisation. Our analysis shows that while most front-end functions exhibit common call graph patterns, a substantial number of rare, divergent graphs occur (Section 5.6). Such rare graphs are easily under-represented, especially under typical trace-sampling policies, resulting in incomplete models and elevated false-positive rates at runtime.
Another assumption of management approaches is that the number of call graphs produced by a deployment is moderate, making such modelling tractable (yang_transparently_2021; panahandeh_serviceanomaly_2024; wang_workflow-aware_2020; meng_detecting_2021). In contrast, we observed 79 million unique call graphs and an average of 44 million call graphs per hour. This scope challenges the computational feasibility of approaches that rely on enumerating or clustering call graphs to provide real-time insights into production systems.
The spread of workload across MSs is highly uneven: many auto-scaling and resource-management approaches implicitly assume that most MSs experience substantial and predictable demand, justifying uniformly sophisticated workload-prediction models (luo_power_2022; choi_phpa_2022; abdullah_burst_2022; nguyen_graph_2022; cheng_proscale_2023). Our findings (Section 5.2), consistent with prior observations (luo_characterizing_2021; luo_power_2022), reveal a strongly long-tailed distribution of workload, with a small fraction of MSs handling the bulk of requests. This suggests that the greatest gains in resource efficiency and service level agreement compliance will come from improving prediction for high-demand MSs, whereas the majority may be adequately handled by simpler heuristic or rule-based scaling.
MS instances are deployed across a large number of nodes and can be co-located with a large number of different MS instances: many management approaches leverage distributed tracing data to capture the direct dependencies between MSs (liu_microhecl_2021; yang_transparently_2021; panahandeh_serviceanomaly_2024; yu_microrank_2021; yu_tracerank_2023; gan_sage_2021; gan_sleuth_2024), implicitly assuming that faults only propagate along these edges. Our analysis hows that MS instances are deployed across a large number of nodes and are often co-located with many other MSs (5.1). Such co-location enables indirect fault propagation through shared resources even in the absence of direct calls (ma_automap_2020). This finding challenges the adequacy of models that rely solely on direct call relationships for reasoning about fault propagation in large-scale MSAs.
6. Dynamicity
MSAs are inherently dynamic. Whilst dependencies are fixed in underlying code, at runtime, calls between MSs are transient, only occurring based on a front-end request. Furthermore, the deployment of MS instances is elastic, with instances being scaled up and down in response to changes in demand and performance. Furthermore, the architecture is prone to change as MSs are added, removed, and updated to evolve the functionality of the system (newman_building_2021). In this section, we explore how this dynamicity manifests in the dataset.